OpenStudy (anonymous):

Differentiate the function y = f(x) = (xcubed − 1)(xcubed + 1).

OpenStudy (kira_yamato):

Hope this helps...

OpenStudy (anonymous):

kira is correct, or you can mult them first y = f(x) = (xcubed − 1)(xcubed + 1). y=(x^3- 1)(x^3 +1) y=x^6 -1 differentiating y '=6x^5

OpenStudy (anonymous):

thnx

OpenStudy (kira_yamato):

No prob...

OpenStudy (anonymous):

ok have fun solving..lol

OpenStudy (kira_yamato):

hahaha.... Thx :P

OpenStudy (anonymous):

thank you. i have a practice test with out an answer key and wanted to make sure i was doing it right